the length of a rectangular garden is 1 ft less than 3 times its width. If the perimeter of the garden is 46 ft, what is the dimensions of the garden
Answer by minerva(22)

breadth=x
length=3x-1
perimeter=2(l+b)=46
46=2(3x-1+x)
46=2(4x-1)
46=8x-2
46+2=8x
48/8=x
breadth = x=6 feet
so length=6 X 3 - 1
length=17
length=17 feet